# A DSMC method for the space homogeneous multispecies Landau equation

### v1

@@ 1 Introduction
Plasmas are complex systems composed by various charged particles, typically consisting of electrons and multiple types of ions. These charged species interact with one another through long-range Coulomb forces. These interactions are fundamental to several key processes, such as the exchange of energy and momentum, the relaxation towards thermodynamic equilibrium, and the emergence of collective behaviours across different spatial and temporal scales. Kinetic theory offers a solid framework for describing these mechanisms by modelling the evolution of each species through a distribution function in phase space. In this context, the Landau–Fokker–Planck equation serves as an effective representation of collisional plasmas, derived as the small-angle collision limit of the Boltzmann equation. By extending this framework to account for multiple interacting species, we obtain the multispecies Landau equation, which includes cross-collisional effects that drive the system toward global thermal equilibrium.
@@
From a computational viewpoint, the numerical solution of the multispecies Landau equation remains challenging due to the nonlocal structure of its collision operator and the need to preserve fundamental physical properties such as mass, momentum, energy conservation and entropy dissipation. Moreover, the mass ratio makes the problem stiff, see e.g. [carrillo2024particle], especially when trying to capture the proton-electron mass ratio $m_{p}/m_{e}\approx 1836$.
@@
To tackle the inherent challenges of numerically solving the Landau equation, both in multispecies and single-species scenarios, various stochastic and Monte Carlo methods have been developed. Among these, the Direct Simulation Monte Carlo (DSMC) methods have gained significant prominence, offering a stochastic representation of binary interactions [bobylev2000, dimarco2010, medaglia2024JCP, medaglia2025dsmc]. In addition to DSMC, alternative formulations based on the drift–diffusion (Fokker–Planck) description of the Landau operator have been proposed [dimits2013higher, manheimer1997langevin]. Recent multi-level extensions have been designed to improve variance reduction in these methods [rosin2014multilevel]. Although these approaches are physically consistent and relatively straightforward to implement, they face limitations due to statistical noise and slow convergence rates. Nevertheless, their computational efficiency and scalability often make them the preferred choice for large-scale plasma applications, as seen in the NESO [threlfall2023software] and XGC1 [chang2008spontaneous, ku2009full] codes. For further applications related to the multispecies Boltzmann equation and Asymptotic Preserving (AP) schemes, see [zhang2016asymptotic, jin2010micro, jin2013bgk].
@@
A distinct line of research has focused on developing deterministic particle methods to more accurately reproduce the dynamics of the Landau operator. A significant advancement in this area is the structure-preserving particle method introduced in [carrillo2020] and extended to the multispecies scenario in [carrillo2024particle, zonta2022multispecies]. This method utilizes the variational gradient-flow formulation of the Landau equation to create a conservative and entropy-dissipating scheme. The method’s computational cost can be considerably lowered using random-batch strategies as discussed in [carrillo2022_randombatch]. Additionally, it has been adapted for various contexts, including spatially inhomogeneous scenarios [bailo2024], uncertainty quantification [bailo2025uncertainty]. The theoretical foundations of this method are underpinned by rigorous convergence analyses presented in [carrillo2022boltzmann, carrillo2023convergence].
@@
In addition to particle-based models, several deterministic approaches based on grid discretisation of velocity space have been developed. These include entropy-consistent methods [buet1998conservative, crouseilles2004numerical, degond1994entropy], finite-difference and finite-volume techniques [taitano2015mass, taitano2016adaptive], as well as fast algorithms utilizing multipole expansions [lapenta2017, lemou2004multipole] or FFT-based spectral methods [dimarco2015, pareschi2000fast, FiPa2002]. While these Eulerian formulations typically achieve higher numerical precision and maintain the essential invariants of Landau dynamics, they can become cumbersome in high-dimensional spaces or complex geometries and are limited by the boundaries of the velocity domain. For a comprehensive review of these advancements, see [dimarco2014numerical]. Very recent findings on inelastic multispecies Boltzmann equation can be found in [rey2025boltzmann].
@@
This study enhances the existing research by expanding the DSMC methodology into the multispecies Landau framework. A key aspect of our approach is the relationship between the Landau and Boltzmann operators in the limit of grazing collisions, which involves small deflection angles. This limit, thoroughly analysed in previous work [bobylev2000], serves as a formal link between stochastic Boltzmann solvers and Landau-Fokker–Planck dynamics [nanbu1997theory, nanbu1997]. Building on this connection, we adopt a first-order approximation of the Boltzmann operator and introduce a regularized scattering kernel, following the ideas presented in [medaglia2024JCP, medaglia2025dsmc]. This approach eliminates the need for iterative solvers while ensuring the physical invariants of the Landau operator are maintained. It is important to note that, while this paper focuses specifically on the spatially homogeneous multispecies Landau equation, the resulting collisional dynamics can be integrated with standard Particle-In-Cell (PIC) solvers for the Vlasov–Maxwell system [dimarco2010, medaglia2023JCP]. A DSMC–PIC coupling has already been developed in the single-species context [medaglia2025dsmc], and the particle-based, mesh-free nature of our multispecies method makes it well-suited for extending this approach to fully multispecies, spatially inhomogeneous simulations.
@@ 2 From the multispecies Boltzmann equation to themultispecies Landau equation
In this section, we introduce the necessary notation and recall the approximation of the multispecies Boltzmann collision operator that leads to the multispecies Landau equation in the grazing-collision limit. The derivation follows the classical approach introduced by Bobylev and Nanbu in their seminal work [bobylev2000], and related developments for the single-species case in [dimarco2010, medaglia2024JCP, medaglia2025dsmc], where the sampling of the angles from a regular distribution has been deeply studied. This approximation is necessary for the construction of the DSMC method that will be presented later on in Section 3.
@@
that depends on the specific form of $\sigma_{\alpha\beta}$. In particular, by recalling the Rutherford formula and the usual cut-off of the angles (see, e.g., [bobylev2000]), for Coulomb collisions we have
@@
It is usual to consider the typical relative speed, and thus the Coulomb logarithm, as a constant in numerical simulations, see e.g. [bobylev2000, caflisch2008hybrid, dimits2013higher]. In this work, we will adopt the same assumption, although a time dependent Coulomb logarithm can be used as well. In the case of Maxwellian interactions we simply have
@@ 2.1 First order approximation
We are now interested in the explicit action of the operator $\exp(\varepsilon_{\alpha}\hat{J}_{\alpha\beta})$ on the angular variable of $F_{\alpha\beta}$ into the gain operator. Generalizing the results of [bobylev2000] (and those of [dimarco2010, medaglia2024JCP, medaglia2025dsmc]) to the multispecies case, we obtain
@@
In practice, sample the components $(\theta,\phi)$ of the solid angle from $D_{\alpha\beta}(\cdot)$ can be numerically challenging. To address this, simpler collisional kernels $D_{\alpha\beta,*}(\cdot)$ may be employed. For a general discussion on the properties of such simpler kernels, we refer to Section VI of [bobylev2000] and Section 2.1 of [medaglia2025dsmc]. Here, we choose
@@
and $\delta(\cdot)$ Dirac delta, following the proposal of [medaglia2024JCP]. The sampling of the components $(\theta,\phi)$ of the solid angle prevents the use of iterative solvers by simply fixing
@@
As discussed in [medaglia2024JCP, medaglia2025dsmc], different formulations of the approximate collision kernel derived in the grazing-collision limit exhibit similar numerical behaviour when applied to both the space-homogeneous and space-inhomogeneous Landau equations for a single-species. This consistency has been verified against standard benchmarks and confirmed through the numerical tests presented in previous works. However, it has been noted that the particular regularisation chosen for the kernel can significantly affect the stability and smoothness of the resulting approximation, especially when the method is extended to uncertainty quantification (UQ) settings. In this context, smooth regularisations of the collision kernel offer enhanced robustness and are therefore preferred for practical implementations. The same considerations apply to the multispecies case.
@@ 4 Numerical results
In all the tests, we will consider the case of two-species, ideally electrons $\alpha$ and ions $\beta$, with two different mass ratios: $m_{\beta}/m_{\alpha}=2$ as a proof of concept, and $m_{\beta}/m_{\alpha}=1836$ which is approximately the real proton-electron mass ratio. We fix $n_{\alpha}=n_{\beta}=1$, $e_{\alpha}=e_{\beta}=1$, $\varepsilon_{0}=1$, $k_{B}=1$, and $\log\Lambda$ is fixed as (8). We also use, following [bobylev2000]
@@
but any other reasonable choice is allowable. The distribution function is always reconstructed with histograms, with $N_{v}=100$ bins per dimension in the domain $L_{\alpha},+L_{\alpha}]\times[-L_{\alpha},+L_{\alpha}]\times[-L_{\alpha},+L_{\alpha}]$ and $L_{\beta},+L_{\beta}]\times[-L_{\beta},+L_{\beta}]\times[-L_{\beta},+L_{\beta}]$ with $L_{\beta}=L_{\alpha}\sqrt{m_{\alpha}/m_{\beta}}$. This proportionality is taken from [carrillo2024particle], however in our case it is not important, because it does not affect the dynamics, since DSMC methods are mesh free. It is simply a choice for the domain of the histogram reconstruction.
@@ 4.1 Test 1: BKW solution for Maxwellian interactions
In dimension $d=3$, the exact-in-time BKW solution to the multispecies Landau equation (see Appendix of [carrillo2024particle]) with Maxwellian collisions reads
@@ 4.2 Test 2: Relaxation towards the equilibrium with Coulomb interactions
Figure 4: Test 2: Relaxation to the equilibrium $m_{\beta}/m_{\alpha}=2$. Time evolution of the temperature (top-left panel), the $x$-components of the velocities (top-right panel), the $y$-component of the velocities (bottom-left panel), and the $z$-component of the velocities (bottom-right panel), for the system with mass ratio $m_{\beta}/m_{\alpha}=2$. In all the panels, the solid red line is the species $\alpha$, the solid blue line the species $\beta$, and the solid black line represents the total quantities, that are conserved. The number of particles is $N=10^{7}$.
Figure 5: Test 2: Relaxation to the equilibrium $m_{\beta}/m_{\alpha}=1836$. Time evolution of the temperature (top-left panel), the $x$-components of the velocities (top-right panel), the $y$-component of the velocities (bottom-left panel), and the $z$-component of the velocities (bottom-right panel), for the system with mass ratio $m_{\beta}/m_{\alpha}=1836$. In all the panels, the red line (solid or star-solid) is the species $\alpha$, the blue line (solid or circle-solid) the species $\beta$, and the solid black line represents the total quantities, that are conserved. In the top-left panel, we display also the detail for small times of the electron temperature $T_{\alpha}(t)$ to highlight its fast evolution toward the maximum before converging towards the equilibrium. The number of particles is $N=10^{7}$.
@@ Conclusions
We have presented a Direct Simulation Monte Carlo (DSMC) method for the spatially homogeneous multispecies Landau equation, built from a first-order grazing-collision approximation of the Boltzmann operator and a regularised angular kernel. The algorithm is conservative at the binary-interaction level, scalable to high-dimensional velocity spaces, and able to treat realistic mass ratios, demonstrated here up to the physical proton–electron value $m_{\beta}/m_{\alpha}\approx 1836$. Numerical tests (BKW and Coulomb relaxation) confirm the method’s ability to reproduce benchmark behaviour, capture rapid inter-species relaxation induced by large mass ratios, and preserve global invariants, namely mass, momentum, and energy. Being mesh-free and particle-based, the scheme integrates naturally with PIC solvers and is well suited for extensions to the inhomogenous scenario. Indeed, future work will target the extension to the spatially inhomogeneous Vlasov–Maxwell-Landau couplings to enable fully kinetic, multispecies plasma simulations with realistic mass ratios.

### v2

@@ 1 Introduction
Plasmas are complex systems composed by various charged particles, typically consisting of electrons and multiple types of ions [3, 14]. These charged species interact with one another through long-range Coulomb forces. These interactions are fundamental to several key processes, such as the exchange of energy and momentum, the relaxation towards thermodynamic equilibrium, and the emergence of collective behaviours across different spatial and temporal scales. Kinetic theory offers a solid framework for describing these mechanisms by modelling the evolution of each species through a distribution function in phase space. In this context, the Landau–Fokker–Planck equation serves as an effective representation of collisional plasmas, derived as the small-angle collision limit of the Boltzmann equation [16]. By extending this framework to account for multiple interacting species, we obtain the multispecies Landau equation, which includes cross-collisional effects that drive the system toward global thermal equilibrium.
@@
From a computational viewpoint, the numerical solution of the multispecies Landau equation remains challenging due to the nonlocal structure of its collision operator and the need to preserve fundamental physical properties such as mass, momentum, energy conservation and entropy dissipation. Moreover, the mass ratio makes the problem stiff, see e.g. [11], especially when trying to capture the proton-electron mass ratio $m_{p}/m_{e}\approx 1836$.
@@
To tackle the inherent challenges of numerically solving the Landau equation, both in multispecies and single-species scenarios, various stochastic and Monte Carlo methods have been developed. Among these, the Direct Simulation Monte Carlo (DSMC) methods have gained significant prominence, offering a stochastic representation of binary interactions [4, 18, 31, 32, 29]. In addition to DSMC, alternative formulations based on the drift–diffusion (Fokker–Planck) description of the Landau operator have been proposed [21, 28]. Recent multi-level extensions have been designed to improve variance reduction in these methods [37]. Although these approaches are physically consistent and relatively straightforward to implement, they face limitations due to statistical noise and slow convergence rates. Nevertheless, their computational efficiency and scalability often make them the preferred choice for large-scale plasma applications, as seen in the NESO [40] and XGC1 [13, 25] codes. For further applications related to the multispecies Boltzmann equation and Asymptotic Preserving (AP) schemes, see [41, 24, 23].
@@
A distinct line of research has focused on developing deterministic particle methods to more accurately reproduce the dynamics of the Landau operator. A significant advancement in this area is the structure-preserving particle method introduced in [12] and extended to the multispecies scenario in [11, 42]. This method utilizes the variational gradient-flow formulation of the Landau equation to create a conservative and entropy-dissipating scheme. The method’s computational cost can be considerably lowered using random-batch strategies as discussed in [7]. Additionally, it has been adapted for various contexts, including spatially inhomogeneous scenarios [1], uncertainty quantification [2]. The theoretical foundations of this method are underpinned by rigorous convergence analyses presented in [8, 9, 10].
@@
In addition to particle-based models, several deterministic approaches based on grid discretisation of velocity space have been developed. These include entropy-consistent methods [5, 15, 17], finite-difference and finite-volume techniques [38, 39], as well as fast algorithms utilizing multipole expansions [26, 27] or FFT-based spectral methods [19, 35, 22]. While these Eulerian formulations typically achieve higher numerical precision and maintain the essential invariants of Landau dynamics, they can become cumbersome in high-dimensional spaces or complex geometries and are limited by the boundaries of the velocity domain. For a comprehensive review of these advancements, see [20]. Very recent findings on inelastic multispecies Boltzmann equation can be found in [36].
@@
This work extends the existing research by adapting the single-species DSMC methodology to the multispecies Landau framework. A key aspect of our approach is the relationship between the Landau and Boltzmann operators in the limit of grazing collisions, which involves small deflection angles. This limit, thoroughly analysed in the seminal work [4], serves as a formal link between stochastic Boltzmann solvers and Landau-Fokker–Planck dynamics [33, 34]. Building on this connection, we adopt a first-order approximation of the Boltzmann operator and introduce a regularized scattering kernel, following the ideas presented in [31, 32]. This approach eliminates the need for iterative solvers while ensuring the physical invariants of the Landau operator are preserved. In the single-species case, this reformulation allows us to obtain a Boltzmann-type equation for Maxwellian molecules, thus eliminating acceptance-rejection processes on collisions, since the physics of interaction is embodied in the sampling of collision angles. In the multispecies case, the mass ratios affect the collision frequencies, making it impossible to treat the case in the same way as for a single species. It is therefore necessary to reformulate the problem so that, on average, the number of interacting particles per time step is consistent with the physics of the problem. This aspect, together with the implementation of the method itself, represents the core of this work. It is also important to note that, while this paper focuses specifically on the spatially homogeneous multispecies Landau equation, the resulting collisional dynamics can be integrated with standard Particle-In-Cell (PIC) solvers for the Vlasov–Maxwell system [18, 30]. A DSMC–PIC coupling has already been developed in the single-species context [32], and the particle-based, mesh-free nature of our multispecies method makes it well-suited for extending this approach to fully multispecies, spatially inhomogeneous simulations.
@@ 2 From the multispecies Boltzmann equation to themultispecies Landau equation
In this section, we introduce the necessary notation and recall the approximation of the multispecies Boltzmann collision operator that leads to the multispecies Landau equation in the grazing-collision limit. The derivation follows the classical approach introduced by Bobylev and Nanbu in their seminal work [4], and related developments for the single-species case in [18, 31, 32], where the sampling of the angles from a regular distribution has been deeply studied. This approximation is necessary for the construction of the DSMC method that will be presented later on in Section 3.
@@
that depends on the specific form of $\sigma_{\alpha\beta}$. In particular, by recalling the Rutherford formula and the usual cut-off of the angles (see, e.g., [4]), for Coulomb collisions we have
@@
It is usual to consider the typical relative speed, and thus the Coulomb logarithm, as a constant in numerical simulations, see e.g. [4, 6, 21]. In this work, we will adopt the same assumption, although a time dependent Coulomb logarithm can be used as well. In the case of Maxwellian interactions we simply have
@@ 2.1 First order approximation
We are now interested in the explicit action of the operator $\exp(\varepsilon_{\alpha}\hat{J}_{\alpha\beta})$ on the angular variable of $F_{\alpha\beta}$ into the gain operator. Generalizing the results of [4] (and those of [18, 31, 32]) to the multispecies case, we obtain
@@
In practice, sample the components $(\theta,\phi)$ of the solid angle from $D_{\alpha\beta}(\cdot)$ can be numerically challenging. To address this, simpler collisional kernels $D_{\alpha\beta,*}(\cdot)$ may be employed. For a general discussion on the properties of such simpler kernels, we refer to Section VI of [4] and Section 2.1 of [32]. Here, we choose
@@
and $\delta(\cdot)$ Dirac delta, following the proposal of [31]. The sampling of the components $(\theta,\phi)$ of the solid angle prevents the use of iterative solvers by simply fixing
@@
As discussed in [31, 32], different formulations of the approximate collision kernel derived in the grazing-collision limit exhibit similar numerical behaviour when applied to both the space-homogeneous and space-inhomogeneous Landau equations for a single-species. This consistency has been verified against standard benchmarks and confirmed through the numerical tests presented in previous works. However, it has been noted that the particular regularisation chosen for the kernel can significantly affect the stability and smoothness of the resulting approximation, especially when the method is extended to uncertainty quantification (UQ) settings. In this context, smooth regularisations of the collision kernel offer enhanced robustness and are therefore preferred for practical implementations. The same considerations apply to the multispecies case.
@@ 3 A Direct Simulation Monte Carlo method
Remark 3.
We have presented the DSMC algorithm in the Nanbu–Babovsky formulation. This choice is motivated by the fact that the Nanbu–Babovsky scheme preserves total mass, momentum, and energy at the discrete level, which is essential for the practical simulations. Moreover, once the collision pairs are selected, the scheme allows for an efficient parallelization of the binary collision step. Of course, other DSMC formulations may be extended to the multispecies case in the same spirit of the one proposed in this work.
Remark 4.
For a multispecies Boltzmann model with $N$ particles per species, the computational cost of a DSMC method scales linearly with the total number of particles. At each time step, the collision stage accounts for both intraspecies and interspecies interactions, whose expected total cost is $O(N)$ per species. As a result, the overall computational complexity per time step remains linear in the total number of particles, even in the multispecies setting.
@@ 4 Numerical results
In all the tests, we will consider the case of two-species, ideally electrons $\alpha$ and ions $\beta$, with two different mass ratios: $m_{\beta}/m_{\alpha}=2$ as a proof of concept, and $m_{\beta}/m_{\alpha}=1836$ which is approximately the real proton-electron mass ratio. We fix $n_{\alpha}=n_{\beta}=1$, $e_{\alpha}=e_{\beta}=1$, $\varepsilon_{0}=1$, $k_{B}=1$, and $\log\Lambda$ is fixed as (8). We also use, following [4]
@@
and analogously for $\pi_{\alpha\alpha}$, $\pi_{\beta\alpha}$, and $\pi_{\beta\beta}$, with the indices adjusted accordingly, but any other reasonable choice satisfying the normalization condition is allowable. The distribution function is always reconstructed with histograms, with $N_{v}=100$ bins per dimension in the domain $L_{\alpha},+L_{\alpha}]\times[-L_{\alpha},+L_{\alpha}]\times[-L_{\alpha},+L_{\alpha}]$ and $L_{\beta},+L_{\beta}]\times[-L_{\beta},+L_{\beta}]\times[-L_{\beta},+L_{\beta}]$ with $L_{\beta}=L_{\alpha}\sqrt{m_{\alpha}/m_{\beta}}$. This proportionality is taken from [11], however in our case it is not important, because it does not affect the dynamics, since DSMC methods are mesh free. It is simply a choice for the domain of the histogram reconstruction. Finally, we choose in all the following tests $N=5\cdot 10^{7}$ particles, and $L_{\alpha}=5$.
@@ 4.1 Test 1: BKW solution for Maxwellian interactions
In dimension $d=3$, the exact-in-time BKW solution to the multispecies Landau equation (see Appendix of [11]) with Maxwellian collisions reads
@@ 4.2 Test 2: Relaxation towards the equilibrium with Coulomb interactions
Figure 4: Test 2: Relaxation to the equilibrium $m_{\beta}/m_{\alpha}=2$. Time evolution of the temperature (top-left panel), the $x$-components of the velocities (top-right panel), the $y$-component of the velocities (bottom-left panel), and the $z$-component of the velocities (bottom-right panel), for the system with mass ratio $m_{\beta}/m_{\alpha}=2$. In all the panels, the solid red line is the species $\alpha$, the solid blue line the species $\beta$, and the solid black line represents the total quantities, that are conserved. The number of particles is $N=5\cdot 10^{7}$.
Figure 5: Test 2: Relaxation to the equilibrium $m_{\beta}/m_{\alpha}=1836$. Time evolution of the temperature (top-left panel), the $x$-components of the velocities (top-right panel), the $y$-component of the velocities (bottom-left panel), and the $z$-component of the velocities (bottom-right panel), for the system with mass ratio $m_{\beta}/m_{\alpha}=1836$. In all the panels, the red line (solid or star-solid) is the species $\alpha$, the blue line (solid or circle-solid) the species $\beta$, and the solid black line represents the total quantities, that are conserved. In the top-left panel, we display also the detail for small times of the electron temperature $T_{\alpha}(t)$ to highlight its fast evolution toward the maximum before converging towards the equilibrium. The number of particles is $N=5\cdot 10^{7}$.
@@ Conclusions
We have presented a Direct Simulation Monte Carlo (DSMC) method for the spatially homogeneous multispecies Landau equation, built from a first-order grazing-collision approximation of the Boltzmann operator and a regularised angular kernel. The algorithm is conservative at the binary-interaction level, scalable to high-dimensional velocity spaces, and able to treat realistic mass ratios, demonstrated here up to the physical proton–electron value $m_{\beta}/m_{\alpha}\approx 1836$. Numerical tests (BKW and Coulomb relaxation) confirm the method’s ability to reproduce benchmark behaviour, capture rapid inter-species relaxation induced by large mass ratios, and preserve global invariants, namely mass, momentum, and energy. Being mesh-free and particle-based, the scheme integrates naturally with PIC solvers and is well suited for extensions to the inhomogenous scenario. Indeed, future work will target the extension to the spatially inhomogeneous Vlasov–Maxwell-Landau couplings to enable fully kinetic, multispecies plasma simulations with realistic mass ratios. Another interesting future research direction is the study of multispecies DSMC algorithm with time dependent collision frequencies $\pi_{\alpha\beta}$ and Coulomb logarithm, to better reproduce the physics of the system.